OpenStudy (princesssleelee):

Find the non-extraneous solutions of the square root of the quantity x plus 9 minus 5 equals quantity x plus 4. x = −2 x = −9 and x = −2 x = 9 x = −9 and x = −8

OpenStudy (princesssleelee):

\[\sqrt{x+9}-5=x+4\]

OpenStudy (princesssleelee):

I found out that the two x's however equal -2 and -26, if i did the math correctly.

OpenStudy (princesssleelee):

-36, not -26, sorry.

OpenStudy (princesssleelee):

no, it doesnt.

OpenStudy (mathstudent55):

\(\sqrt{x + 9} - 5 = x + 4\) Add 5 to both sides.

OpenStudy (mathstudent55):

Then square both sides.

OpenStudy (solomonzelman):

non-extraneous solution: The solution that works in the original equation. extraneous solution: The solution that is the result of (correct) work/transformations, but that does not work in the initial equation. (definitions, just in case)

OpenStudy (mathstudent55):

\(\sqrt{x + 9} = x + 9\) \((\sqrt{x + 9})^2 = (x + 9)^2\)

OpenStudy (princesssleelee):

I got past that part

OpenStudy (solomonzelman):

can you post the work that you have please?

OpenStudy (mathstudent55):

\(x + 9 = x^2 + 18x + 81\)

OpenStudy (princesssleelee):

then subtract x

OpenStudy (mathstudent55):

\(x^2 + 17x + 72 = 0\)

OpenStudy (princesssleelee):

eventually getting x=-2 and x=-36

OpenStudy (princesssleelee):

I passed all of that, and thats what I got

OpenStudy (mathstudent55):

2 and 36 do not add to 17, so that is incorrect. You need to factor the left side correctly.

OpenStudy (mathstudent55):

\(x^2 + 17x + 72 = 0\) At this point, you can try factoring the left side. What two numbers have a product of 72 and a sum of 17?

OpenStudy (princesssleelee):

Now i am completely lost. I did not add 17

OpenStudy (mathstudent55):

\(x^2 + 17x + 72 = 0\)

OpenStudy (mathstudent55):

Wait a minute. Let's see where your error starts. Did you get the equation I got?

OpenStudy (princesssleelee):

I am going to try and upload a photo of my work from where we were on the same page

OpenStudy (mathstudent55):

Ok

OpenStudy (princesssleelee):

Sorry, it is sloppy and i wasnt planning on sharing lol

OpenStudy (princesssleelee):

I will also share the example that I have in my notes.

OpenStudy (mathstudent55):

Starting with the last line as line 1, go up to line 6. You have \(0 = x^2 + 17x + 72\) You have the correct equation. Ok?

OpenStudy (mathstudent55):

Your work is very clear. No problem there.

OpenStudy (mathstudent55):

Your error is on a line below where you tried to factor the quadratic.

OpenStudy (princesssleelee):

What was the error

OpenStudy (mathstudent55):

You factored like this: \((x + 2)(x + 36) \) If you multiply it out (using FOIL), you get this: \(= x^2 + 36x + 2x + 72\) \(= x^2 + 38x + 72\) As you can see this is not the quadratic you started with.

OpenStudy (princesssleelee):

what did I start with?

OpenStudy (mathstudent55):

You started wit the quadratic you wanted to factor: \(x^2 + 17x + 72\)

OpenStudy (mathstudent55):

Ok. Now we know that your factoring is not correct since it does not multiply out back to the quadratic trinomial you had originally.

OpenStudy (mathstudent55):

Your factoring is of this type: